Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: In CnB configuration; many deferred changes in the system (particularly correlator-based) so expecting impact on science; attempted science restart last night but not quite ready yet (more in Correlator section).
    • News from the ALMA Spectroscopy workshop.
    • Propose no meeting next Tuesday in honor of the upcoming proposal submission deadline and the ongoing move activities.
  • Key dates: * 01 Feb: Proposal deadline for D configuration+ * 03-04 Feb: Dish verification review ($KA)
  • Staff Travel:
  • Visitors:
    • 05 Jan - 06 Apr: John Cannon sabbatical at DSOC
    • New RSRO Scientists: Sergio Dzib, Hendrik Linz, Ashley Zauderer

Commissioning priorities for the week (McMullin, Chandler)

Single maintenance day week:
  • Mon-Tue, Thu-Fri:
    • 0900-1700 Testing/software development time
    • 1800-0900: Science operations
  • Wed: Maintenance day
  • Interrupt status for additional 3-bit testing

Hardware status and updates (Langley, Perley, Rupen)

Software status and updates (Butler)

Commissioning Reports

Calibration: Perley: Cross-observatory flux calibrator survey

  • Followup details: Mostly good stability less than 1% bandpass for most; some indication of attenuators not 'remembering' between band changs leading to 1-2% issues with bandpass; key problem. Other signs of issues with phase stability at X, Ku; switched power testing shows good correction for variations (diurnal in receivers) with the exception of 1-2 antennas.

Spectral Line: Goss/Chomiuk: Bandpass stability monitoring

  • Full report postponed to Laura's visit in February (1 Feb meeting). Early results posted at ECSV-107; integrating information from the 30 hour run.

Polarization: Momjian/Sowinski: Zeeman testing/V stokes issues

  • Status report; see: ECSV-105.

Calibration: Hunter/Brogan/McMullin: Residual delays in observations

  • Status/discussion: no differences between wide-band and narrow-band observations. Overall seeing delays of ~2-3 ns which is currently within the operating specification regime. However, Todd has noted that ALMA's spec is much smaller (1 picosecond) and errors on the order of our 'standard' will still lead to dynamic range issues if not solved for. There are apparently two different routines available that need to be tested out for CASA reduction: K term calibration in gaincal; multi-band group delay script from Nick (CAS-2618).

Commissioning Group updates and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

RSRO/ECSO Reports

  • Excellent series of talks at the AAS (already discussed at science tea).

Individual reports (updates)

Misc:

Demo Science Status

AOB

  • Action: Joe: Nail down Paul Demorest visit. Proposed last two weeks of Adam's work but this remains unconfirmed.
  • Action: Joe: Push new target to Scheduling group
  • Action: Joe: Finalize report to engineers on EA05 ripple.From Kumar, posted information on the L band ripple; still present.
  • Action: Miriam: Review priorities with Jeff to see if integration of autoflag can be raised for this cycle. Discussion begun.
  • Action: Joe: Post Urvashi's autoflag test information - get feedback for next meeting.Details on using the LOFAR package for autoflag are posted at: https://bugs.nrao.edu/browse/CAS-2676

Misc Actions

  • [Frail/Butler]: Enable restrictions based on proximity to sun; review OST weather trending.
  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.
-- JosephMcMullin - 2011-01-21
Topic revision: r2 - 2011-01-21, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback